OpenStudy (anonymous):

-9 = 5/17n? n= - 153/5 n= - 5/153 n = -17/45 n= -45/17 helpp???

OpenStudy (anonymous):

Cross Multiply!

OpenStudy (anonymous):

-9*17n =5 -153 = 5 n = -5/153